40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12802     Entry time: Mon Feb 6 10:05:28 2017
Author: Steve 
Type: Update 
Category: SUS 
Subject: clamped cables 

The bottom 5 cable connections from Sat-Amp to Whittering Filter at 1X5 were clamped today.

Attachment 1: clamped.jpg  66 kB  | Hide | Hide all
clamped.jpg
ELOG V3.1.3-